OpenStudy (anonymous):

Enter the quadratic equation in standard form in the box below. x^2-5x-5=3x+2

OpenStudy (anonymous):

You need to add like terms.

OpenStudy (anonymous):

Standard form is Ax+By=C

OpenStudy (anonymous):

yah but whats the answer

OpenStudy (anonymous):

\[x ^2 -8x = 7 \] because you need to subtract the 3x and add the 5 to other side

OpenStudy (anonymous):

x^2-8 x-7 http://www.mathsisfun.com/algebra/quadratic-equation.html

OpenStudy (anonymous):

Ok so the new equation states that x^2 -5x -5=3x+2 ..so u set up a new equation by simplifying it and it should look like this : 3x+2=x2 because when. You combine like terms negative 5 and negative 5 cancel add because they add to zero. So once u have the equation 3x+2= x2 you do -2 on both sides of the equation bring down the 3x and u are remained with 3x =-1 the you divide 3 on both sides and u get the answer of -.3